A HIGH RESOLUTION STUDY OF THREE TOP INTERNAL ROTORS: THE MICROWAVE SPECTRA OF TRIMETHYLSILANE,(CH_{3})_{3}SiH, TRIMETHYLSILYLACETYLENE, (CH_{3})_{3}SiC≡CH, and TRIMETHYLSI-LYLDIACETYLENE, (CH_{3})_{3}SiC≡C-C≡CH

\underline{LU KANG}, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The microwave spectra of (CH_{3})_{3}SiH, (CH_{3})_{3}SiC≡CH, and (CH_{3})_{3}SiC≡C-C≡CH, as well as some of their Si isotopomers have been studied with FTMW spectroscopy. (CH_{3})_{3}SiC≡C-C≡CH was generated by passing a jet, consisting of a 1% one to one mixture of (CH_{3})_{3}SiC≡CH and acetylene in an argon carrier gas, through a 900 volt discharge. The internal rotational splitting of (CH_{3})_{3}SiH and (CH_{3})_{3}SiC≡CH have been observed.